# Data-Copilot: Bridging Billions of Data and Humans with Autonomous Workflow

### v7

@@ F.1 Prompts For Data Exploration
F.2 Prompts For Workflow Deployment
@@ F.3 Prompts For Baselines
Appendix G Case Study
Detailed cases of data exploration and workflow deployment are presented in Sections G.1 and G.2.
G.1 Cases for Data Exploration
@@
G.2 Cases for Real-time Request
@@
Figure G2: Data-Copilot deploys workflows to solve users’ prediction request. It invokes three interfaces step by step and generates arguments for each interface.
Figure G3: We have listed some of Data-Copilot’s self-design interfaces, including five categories. Most of the interfaces are a combination of several simple interfaces. For example, the text marked in blue indicates that the interface includes multiple functions.
Figure G4: For complex requests about stock financial data, Data-Copilot deploys a loop workflow to solve user requests and finally outputs images and tables in parallel.
Figure G5: Data-Copilot acquires the P/E data of two stocks in parallel and plots them in a graph.
Figure G6: Data-Copilot also has access to fund data and can query the returns of all funds managed by the fund manager.
Figure G7: Data-Copilot can plot multiple indicators in a single graph by deploying workflows in parallel.
Figure G8: Data-Copilot can provide the latest financial news and company information by deploying the corresponding workflows.
Figure G9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.

### v8

@@ E.3 Prompts For Data Exploration
E.4 Prompts For Workflow Deployment
@@ E.5 Prompts For Baselines
Appendix F Case Study
Detailed cases of data exploration and workflow deployment are presented in Sections F.1 and F.2.
F.1 Cases for Data Exploration
@@
F.2 Cases for Real-time Request
@@
Figure F2: Data-Copilot deploys workflows to solve users’ prediction request. It invokes three interfaces step by step and generates arguments for each interface.
Figure F3: We have listed some of Data-Copilot’s self-design interfaces, including five categories. Most of the interfaces are a combination of several simple interfaces. For example, the text marked in blue indicates that the interface includes multiple functions.
Figure F4: For complex requests about stock financial data, Data-Copilot deploys a loop workflow to solve user requests and finally outputs images and tables in parallel.
Figure F5: Data-Copilot acquires the P/E data of two stocks in parallel and plots them in a graph.
Figure F6: Data-Copilot also has access to fund data and can query the returns of all funds managed by the fund manager.
Figure F7: Data-Copilot can plot multiple indicators in a single graph by deploying workflows in parallel.
Figure F8: Data-Copilot can provide the latest financial news and company information by deploying the corresponding workflows.
Figure F9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.